OpenStudy (anonymous):

Need Help! Will Medal! Which points satisfy this linear-quadratic system of equations? x + y = 1 y = 4x − 3 − x2 @ganeshie8 @hartnn @Directrix @iambatman @dan815 @campbell_st

OpenStudy (campbell_st):

well you may well have 2 points rewrite the 1st equation with y as the subject x + y = 1 so y = ?

OpenStudy (campbell_st):

any thoughts @Firefly007

OpenStudy (campbell_st):

or you can wait and someone will give the answer

OpenStudy (shaik0124):

x+y-x=1-x y=1-x

OpenStudy (anonymous):

(1, 0) and (4, -3) ?

OpenStudy (campbell_st):

ok... so someone has done that for you... so now equate the 1 equations \[1 - x = -3 + 4x - x^2\] collect like terms and have all terms on one side of the equation. what would you get..?

OpenStudy (campbell_st):

so it that your answer or the choices..?

OpenStudy (anonymous):

wait i got it right :) thanks for helping

OpenStudy (anonymous):

:3